OpenStudy (anonymous):

at what values is the expression 4/6-5x+x^2 undefined

OpenStudy (tkhunny):

4/6-5x+x^2 There are no values for x which make this expression undefined. 4/(6-5x+x^2) THIS expression, however, may have two. Factor the denominator and see where that thing is zero (0).

OpenStudy (anonymous):

how would i do that?

OpenStudy (tkhunny):

You've never factored a quadratic equation?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

Then you cannot solve this problem easily. You should have been given this tool before you were given this problem. Well, just figure out where the denominator takes on zero (0). You will have to hunt for the two places where this occurs. Perhaps make a table. I would suggest looking between x = 4 and x = -4. Also, please acknowledge that you know the difference between the two expressions. The one you wrote, which has no such places and the one I corrected, which has two such places.

OpenStudy (anonymous):

so would the answer be it is defined for all values of x

OpenStudy (tkhunny):

For the one you write, which is incorrect, that would be a good answer. For the one you intended, that you did not write, that is not a good answer. Clarify. Did you mean this: \(\dfrac{4}{6 - 5x + x^{2}}\)?

OpenStudy (anonymous):

yes that is the problem i am looking at now

OpenStudy (tkhunny):

Do you understand that this is NOT what you wrote in your original post? You wrote \(\dfrac{4}{6} -5x + x^{2}\)

OpenStudy (anonymous):

Yes I did what do I do next?

OpenStudy (tkhunny):

Asked and answered. Figure out where that denominator is zero (0). Solve: \(6 - 5x + x^{2} = 0\)

OpenStudy (anonymous):

Ok what do I do after that?

OpenStudy (tkhunny):

You interpret the results and answer the question.

OpenStudy (anonymous):

How would i do that I'm still stuck at solving the equation.?

OpenStudy (tkhunny):

Then why do you care what's next? Again, asked and answered. Evaluate the expression for values between x = 4 and x = -4 and see where it takes on the value zero (0). Fill in the table: x = -4 -- Expression = ?? x = -3 -- Expression = ?? x = -2 -- Expression = ?? x = -1 -- Expression = ?? x = 0 -- Expression = ?? x = 1 -- Expression = ?? x = 2 -- Expression = ?? x = 3 -- Expression = ?? x = 4 -- Expression = ??

OpenStudy (anonymous):

can you please just give me the answer?

OpenStudy (tkhunny):

No. I will not do it for you. You have what you need. Show your work and I will be delighted to take a look at your results and conclusions.

OpenStudy (anonymous):